two planes which are 1905 miles apart, fly towards each other. their speed differ by 35 mph.If they pass each other in 3 hours what is the speed of each?
Answer by macston(5194)

r=rate of slower plane; r+35mph=rate of faster plane
1905mi/(r+r+35mph)=3 hrs
1905mi/(2r+35mph)=3hrs Multiply each side by 2r+35mph.
1905mi=3hr(2r+35mph)Divide each side by 3 hr.
635mph=2r+35mph Subtract 35 mph from each side.
600mph=2r Divide each side by 2.
300mph=r ANSWER 1: The rate of the slower plane is 300 mph.
r+35 mph=300mph+35mph=335 mph ANSWER 2: The faster plane traveled at 335 mph.